OpenStudy (anonymous):

Solve: 4x^2+10x=-1 I'd like somebody to walk me through the steps.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

add 1 to both sides 4x^2+10+1=0 Do you know the quadratic formula?

OpenStudy (anonymous):

or just transpose -1 one to the left then get the opposite sign

OpenStudy (anonymous):

bro you got like 4x^2 +10 - 1 = 0 you cant factor so you gotta quadratic forumal that -b (plus or minus) Root(b^2-4AC) all divided by 2A

OpenStudy (anonymous):

that is supposed to be a plus 1 not a minus

OpenStudy (gw2011):

The equation becomes: 4x^2+10x+1=0 You need to use the quadratic formula which is: x=[-b+/-square root(b^2-4ac)]/2a a=4 b=10 c=1 x=[-10+/-square root(10^2-4(4)(1)]/2(4) x=[-10+/-square root(100-16)]/8 x=[-10+/-square root(84)]/8 x=[-10+/-square root(4)*(21)]/8 x=[-10+/-2(square root 21)/8 You can reduce 10, 2 and 8 since each is divisible by 2 and you get: x=[-5+/-(square root 21)]/4=(-5+square root 21)/4 and (-5-square root 21)/4